OpenStudy (anonymous):

find the value of x that makes the equation 1.5x + 6 = 3(2.5x - 6) True.

OpenStudy (mathstudent55):

First, distribute the 3 on the right side. What do you get?

OpenStudy (anonymous):

x=4

OpenStudy (jango_in_dtown):

1.5x + 6 = 3(2.5x - 6) gives 1.5x+6=7.5-18 this gives 6=24 which gives x=4

OpenStudy (jango_in_dtown):

6=24x

OpenStudy (anonymous):

@mathstudent55 is it 4, Also it would be 3(1.5x + 6) and 3(2.5x - 6)

OpenStudy (anonymous):

@jango_IN_DTOWN ty